Solve the following equations. 
z/(z + 15) = 4/9 
Solve the following equations. 
z/(z + 15) = 4/9 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

z/(z+15)=4/9
⇒z=4/9(z+15)
⇒9z=4(z+15)
⇒9z=4z+60
⇒9z-4z=60
⇒5z=60
⇒z=12

Thanks